How do I solve 5x+4=7 and x-3y=9. I have to find the graph
Answer by Alan3354(69443)

How do I solve 5x+4=7 and x-3y=9.
Did you mean 5x+4y=7 ?? I'll assume so.
I have to find the graph
-------------
There are several ways to solve these, I'll use substition.
x-3y=9
x = 3y+9
Sub for x in the 1st eqn
5*(3y+9)+4y = y
15y + 45 + 4y = 7
19y = -38
y = -2
-------
Sub for y into either eqn
x = 3y+9
x = 3*(-2)+9
x = 3